OpenStudy (anonymous):

how do I solve a linear system equation with the two equations being y=-x+4 and x+y=4 by using substitution method

OpenStudy (shoba):

y = -x + 4----1 x+y = 4----2 x + (-x+4)=4 is the equation correct?

OpenStudy (aum):

If you rearrange the given equations a little bit you will notice they are one and the same equation y = -x + 4. So the two equations are dependent and therefore ANY point on the line y = -x + 4 is a solution and there are INFINITE number of solutions.

OpenStudy (aum):

If you follow the substitution method in shoba's reply where you take y from the first equation and put it in the second equation, you get: x - x + 4 = 4 or 4 = 4 which is ALWAYS true no matter what x is. Therefore any value of x is a solution. You just put that x value in y = -x +4 and find the corresponding y value. Thus there are infinite number of solutions.